# Geometric spectral theory of quantum graphs

### v2

@@ 3 Laplacian and Schrödinger operators on metric graphs
3.1 Lemma.
The form $a_{q}$ is a bounded Hermitian form which is $L^{2}$-elliptic on $H^{1}(\Gamma)\times H^{1}(\Gamma)$, that is, there exist $\xi\in\mathbb{R}$ and $\omega>0$ such that
@@
Proof.
Exercise 2.2. ∎
3.2 Proposition.
@@
Proof.
Important exercise (Exercise 2.3). ∎
@@
    $f$ should be continuous at $v$ (this is contained in the condition $f\in H^{1}(\Gamma)\hookrightarrow C(\Gamma)$;
@@
(1) and (2) together are variously called standard, natural, Neumann–Kirchhoff, and continuity–Kirchhoff vertex conditions, and there are probably other names in use as well. (Observe that if $\deg v=1$, then they reduce to just a Neumann condition at the endpoint of the edge.) When $q=0$, $A_{0}$ is often called the Laplacian with standard vertex conditions, a.k.a. the standard or Kirchhoff Laplacian.
3.3 Remark.
@@
4 First observations on the spectrum
@@
This principle is a partial motivation for much of what we will do in Chapter 2, which is one such approach. There are, however, several others, often very well developed, and although it would go beyond the scope of these notes to explore them here, it seems appropriate to mention a few of what are arguably the most important, and give a few references. This is however by no means a complete list, either in terms of the available techniques or in terms of the literature; for that, we refer to some of the many good books and surveys on the subject, including [10, 14, 36].
First, a systematic approach to the secular equation is offered by scattering matrices and the secular determinant, see, e.g., [6, Section 1.4], [10, Section 5], or [14, Section 2.1].
Second, there are powerful, and often exact, trace formulae available for the spectrum of the quantum graph as a whole, and typically based on the periodic orbits (or, put differently, the set of closed paths on the graph), which results, for example, in exact expressions for the spectral measure, or density of states, of the graph in terms of its geometric and topological properties, or as an inverse formula for the Euler characteristic in terms of the spectrum. There is an advanced body of literature studying such formulae, which are often inspired by similar but generally not exact formulae on manifolds; in the context of quantum graphs these probably go back to the seminal paper [34], see also, e.g., [7] or, quite recently, [36, Chapters 8 and 9].
Third, one can use M-functions (known in the literature under a variety of different names, including Titchmarsh–Weyl M-functions and Dirichlet-to-Neumann operators) to study the spectrum in dependence on the vertex coupling conditions (cf. Exercise 2.4), which in turn can be used to analyse the spectrum of even just the standard Laplacian on a graph, see [36, Chapters 17 and 18].
4.4 Example (Secular equation of a star graph).
Cf. [10, Example 2.3]. Consider the $3$-star graph of Example 1.4, but where now $e_{k}$ has length $\ell_{k}>0$, $k=1,2,3$, and suppose $q=0$. We will find the (or, more precisely, a) secular equation for the eigenvalues of $A_{0}$.
@@
Chapter 2 Geometric spectral theory of quantum graphs
1 On domains
Given a bounded domain $\Omega\subset\mathbb{R}^{d}$, $d\geq 2$, with sufficiently smooth boundary, as in Section 2 we denote by
@@ 1 On domains
1.1 Theorem (Faber–Krahn).
@@
Since $\lambda_{1}$ governs the rate of decay of (the $L^{2}$-norms of) solutions to the heat equation (Section 3), this theorem says that, among all objects of given volume, if there is perfect cooling at the boundary, the rate of heat loss (in this sense) is minimised when the object is a ball.
Theorem 1.1 is a starting point of geometric spectral theory, or more specifically shape optimisation in spectral theory, see [24] or [25], which also finds applications in nonlinear PDEs (where the eigenvalues often play the role of a critical or threshold value; an example of this happening on metric graphs is in [16, Section 3]) and related functional inequalities (for example, Theorem 1.1 together with the variational characterisation of the eigenvalues shows that $\lambda_{1}(B)$ is the smallest constant depending only on the volume of the domain for which Friedrichs’ inequality is valid); these, in turn, are often needed in stability and domain perturbation analysis (e.g., [19, 27]). Just like, or even more than, in the case of quantum graphs, only in extremely special cases is it possible to determine the eigenvalues exactly.
@@
1.2 Theorem (Szegő–Weinberger).
@@
The strategy involves symmetrisation (or rearrangement): starting from $0<\psi_{1}\in H^{1}_{0}(\Omega)$ an eigenfunction for $\lambda_{1}(\Omega)$, using tools from geometric measure theory (the coarea formula and the isoperimetric inequality applied to the level sets of $\psi_{1}$) we construct a new function $0<\psi_{1}^{\ast}\in H^{1}_{0}(B)$, a symmetrisation of $\psi_{1}$ such that
@@
2 On graphs: Nicaise’ inequality
@@
We wish to understand how these eigenvalues are related to the geometry (and topology, and metric features) of $\Gamma$, akin to the examples sketched in Section 1. As discussed in Section 4, in general it is essentially impossible to determine the $\mu_{k}$ explicitly (or even to find a convenient representation of the generally extremely complicated transcendental secular equation that characterises them).
@@
2.1 Theorem (Nicaise’ inequality).
@@
This theorem was originally discovered and proved by Nicaise in 1987 [41], rediscovered by Friedlander in 2005 (with a different proof) [23], and then rediscovered again by Kurasov and Naboko in 2014 [38]. The interpretation in terms of diffusion is that, if the graphs are perfectly insulated, so that no heat can escape, then, among all graphs of given total length, the interval exhibits the slowest convergence (in the usual $L^{2}$-sense) to equilibrium $=$ uniform heat distribution. Actually, there is a corresponding result for discrete graph Laplacians which is even older, going back to the 1970s [22], see also [11, Section 2].
@@
2.2 Theorem (Nicaise, Dirichlet case).
@@
[…]
Proof.
@@
secondly, if $y_{t}\in[0,L]$ is the unique point such that $\psi_{1}^{\ast}(y_{t})=t$, i.e., $\{y_{t}\}=S_{t}^{\ast}=\partial U_{t}^{\ast}$, then the relation $|U_{t}|=|U_{t}^{\ast}|$ implies
@@
where the first line is the coarea formula, the second follows from the first observation, the third from the second observation, and the final line is another application of the coarea formula. ∎
@@
2.5 Example.
If $\Gamma=I=[0,L]$ is just a bounded interval, and $\psi_{k}\sim\lambda_{k}$ is a $k$th eigenfunction of $A_{q}$ on $I$ ($q\in L^{\infty}(I)$), then by Sturm–Liouville theory, $\psi_{k}$ has exactly $k-1$ zeros in the interior of $I$, and thus exactly $k$ nodal domains $\Omega_{1},\ldots,\Omega$.
@@
[…]
@@
Proof.
By assumption $\psi_{k}$ satisfies the eigenvalue equation $-\psi_{k}=\mu_{k}\psi_{k}$ strongly (indeed pointwise) in $\Gamma$, and hence in $\Omega$. But it also satisfies all vertex conditions in $\Omega$ by construction. Thus it is equal to some eigenfunction on $\Omega$; in particular, $\mu_{k}=\lambda_{j}(\Omega)$ for some $k\geq 1$.
Now since $\psi_{k}$ does not change sign in $\Omega$, and $\lambda_{1}(\Omega)$ is the only eigenvalue on $\Omega$ with a non-sign-changing eigenfunction, we must have $j=1$. ∎
First proof of Theorem 2.1.
Fix an eigenfunction $\psi_{2}\sim\mu_{2}(\Gamma)$, then $\psi_{2}$ has (at least) two nodal domains $\Omega^{+}$, $\Omega^{-}$; at least one of them, say $\Omega^{+}$, has total length $\leq L/2$. Since $\partial\Omega^{+}\neq\emptyset$, i.e., $\Omega^{+}$ has at least one Dirichlet vertex, by Lemma 2.7 and Theorem 2.2,
@@
Rough idea of proof.
If $\psi_{k}$ has $k$ nodal domains, then at least one has length $\leq L/k$; apply Theorem 2.2. Otherwise, there exists a linear combination of $\psi_{1},\ldots,\psi_{k}$ which does have $k$ nodal domains. Apply the symmetrisation technique to one of these. ∎
2.9 Remark.
@@
On the other hand, the values in Friedlander’s theorem diverge sharply from the values of $\mu_{k}(\Gamma)$ given by Weyl’s law for any fixed graph $\Gamma$. This is starkly at odds with Pólya’s conjecture (see, e.g., [18]) for the Dirichlet and Neumann Laplacian eigenvalues, which posits that the first term in the Weyl asymptotics is always a lower bound for the Dirichlet eigenvalues and an upper bound for the Neumann eigenvalues, on any domain. It seems likely that this is related to the failure of the unique continuation principle mentioned in Section 4 and to the nodal domain counts explored in [26].
3 Surgery
2nd proof of Theorem 2.1: the “doubling trick” (Nicaise/Kurasov–Naboko)
@@ 3 Surgery
[…]
Duplicate the eigenfunction $\psi_{2}\in H^{1}(\Gamma)$ on each edge to create a new function $\tilde{\psi}_{2}\in H^{1}(\Gamma_{2})$; then by construction
@@
3.1 Theorem (Euler, 1736, “Bridges of Königsberg”).
@@
$\mu_{2}(\Gamma)\geq\mu_{2}(\Gamma_{2})=\inf_{\begin{subarray}{c}0\neq u\in H^{1}(\Gamma_{2})\\ \int_{\Gamma_{2}}u\,\textrm{d}x=0\end{subarray}}\frac{\int_{\Gamma_{2}}|u^{\prime}|^{2}\,\textrm{d}x}{\int_{\Gamma_{2}}|u|^{2}\,\textrm{d}x}=\mu_{2}(\Gamma_{2})\geq\inf_{\begin{subarray}{c}0\neq u\in H^{1}(C)\\ \int_{C}u\,\textrm{d}x=0\end{subarray}}\frac{\int_{C}|u^{\prime}|^{2}\,\textrm{d}x}{\int_{C}|u|^{2}\,\textrm{d}x}=\mu_{2}(C)=\mu_{2}(C).$
@@
Actually, this inequality holds for a more general class of graphs, known as doubly connected graphs (Theorem of Band–Lévy, [8, Theorem 2.1(2)]). Intuitively, graphs which are more connected, i.e., have more non-overlapping paths between any pairs of points, necessarily have larger $\mu_{2}$ and thus faster convergence of diffusion processes to equilibrium. Higher graph connectivity and its effect on the eigenvalues is also explored in [11]. This closely parallels results for discrete graph Laplacians which have been known for many decades [22, 39], where the discrete counterpart of $\mu_{2}$ is even called the algebraic connectivity of the graph.
@@
See also Exercise 3.2 or [12, Section 3]. The process of cutting/gluing is more complicated if other vertex conditions are imposed, since then, unlike with the case of standard conditions, there may be no natural conditions the newly created vertices should satisfy.
Proof.
@@
(1) now follows from the min-max principle (Theorem 1.5/Remark 4.1). ∎
3.5 Corollary (Weyl asymptotics).
Let $\Gamma$ be a compact metric graph with total length $|\Gamma|=L$. Then
@@
3.6 Theorem.
(Cf. [30, Theorem 4.2].) Let $\Gamma$ be a compact metric graph with total length $L>0$ and $m$ edges. Then
@@
3.8 Lemma.
Suppose $\widetilde{\Gamma}$ is formed from $\Gamma$ by deleting a pendant subgraph $\Gamma^{\prime}\subset\Gamma$ (equivalently, $\Gamma$ is formed from $\widetilde{\Gamma}$ by gluing a new graph Γsuperscriptnormal-$\Gamma^{\prime}$ to $\widetilde{\Gamma}$ at a single vertex). Suppose standard conditions are imposed at all vertices of all graphs in question. Then
@@
for all $k\geq 1$. ∎
This lemma can also be proved using a test function argument, see [43, Proposition 3.1]. Its sharpest form can be found in [12, Theorem 3.10]. It is intuitively clear that attaching a pendant subgraph (basically, an extra dead end) should slow the rate of diffusion. The picture is far more complicated if the subgraph being attached or deleted is not a pendant, but rather attached at two or more points, as the following simple example shows.
3.9 Example.
@@
It is left as another exercise to check that the second eigenvalue of an equilateral $m$-flower graph is in fact $\frac{\pi^{2}m^{2}}{L^{2}}$. ∎
4 Advanced surgery

### v3

@@ 3.3 Laplacian and Schrödinger operators on metric graphs
Lemma 3.20.
The form $a_{q}$ is a bounded Hermitian form which is $L^{2}$-elliptic on the space $H^{1}(\Gamma)\times H^{1}(\Gamma)$, that is, there exist $\xi\in\mathbb{R}$ and $\omega>0$ such that
@@
Proof 3.21.
Exercise 2.2.
Proposition 3.22.
@@
Proof 3.23.
Important exercise (Exercise 2.3).
@@
    $f$ should be continuous at $v$ (this is contained in the condition $f\in H^{1}(\Gamma)\hookrightarrow C(\Gamma)$);
@@
Conditions (1) and (2) together are variously called standard, natural, Neumann–Kirchhoff, and continuity–Kirchhoff vertex conditions, and there are probably other names in use as well. (Observe that if $\deg v=1$, then they reduce to just a Neumann condition at the endpoint of the edge.) When $q=0$, $A_{0}$ is often called the Laplacian with standard vertex conditions, a.k.a. the standard or Kirchhoff Laplacian.
Remark 3.24.
@@
3.4 First observations on the spectrum
@@
This principle is a partial motivation for much of what we will do in Chapter 4, which is one such approach. There are, however, several others, often very well developed, and although it would go beyond the scope of these notes to explore them here, it seems appropriate to mention a few of what are arguably the most important, and give a few references. This is however by no means a complete list, either in terms of the available techniques or in terms of the literature; for that, we refer to some of the many good books and surveys on the subject, including [11, 15, 37].
First, a systematic approach to the secular equation is offered by scattering matrices and the secular determinant, see, e.g. [6, Section 1.4], [11, Section 5], or [15, Section 2.1].
Second, there are powerful, and often exact, trace formulae available for the spectrum of the quantum graph as a whole, and typically based on the periodic orbits (or, put differently, the set of closed paths on the graph), which results, for example, in exact expressions for the spectral measure, or density of states, of the graph in terms of its geometric and topological properties, or as an inverse formula for the Euler characteristic in terms of the spectrum. There is an advanced body of literature studying such formulae, which are often inspired by similar but generally not exact formulae on manifolds; in the context of quantum graphs these probably go back to the seminal paper [35], see also, e.g. [7] or, quite recently, [37, Chapters 8 and 9].
Third, one can use M-functions (known in the literature under a variety of different names, including Titchmarsh–Weyl M-functions and Dirichlet-to-Neumann operators) to study the spectrum in dependence on the vertex coupling conditions (cf. Exercise 2.4), which in turn can be used to analyse the spectrum of even just the standard Laplacian on a graph, see [37, Chapters 17 and 18].
Example 3.30 (Secular equation of a star graph).
Cf. [11, Example 2.3]. Consider the $3$-star graph of Example 3.2, but where now $e_{k}$ has length $\ell_{k}>0$, $k=1,2,3$, and suppose $q=0$. We will find the (or, more precisely, a) secular equation for the eigenvalues of $A_{0}$.
@@
4 Geometric spectral theory of quantum graphs
4.1 On domains
Given a bounded domain $\Omega\subset\mathbb{R}^{d}$, $d\geq 2$, with sufficiently smooth boundary, as in Section 2.2 we denote by
@@ 4.1 On domains
Theorem 4.1 (Faber–Krahn).
@@
Since $\lambda_{1}$ governs the rate of decay of (the $L^{2}$-norms of) solutions to the heat equation (Section 2.3), this theorem says that, among all objects of given volume, if there is perfect cooling at the boundary, the rate of heat loss (in this sense) is minimised when the object is a ball.
Theorem 4.1 is a starting point of geometric spectral theory, or more specifically shape optimisation in spectral theory, see [25] or [26], which also finds applications in nonlinear PDEs (where the eigenvalues often play the role of a critical or threshold value; an example of this happening on metric graphs is in [17, Section 3]) and related functional inequalities (for example, Theorem 4.1 together with the variational characterisation of the eigenvalues shows that $\lambda_{1}(B)$ is the smallest constant depending only on the volume of the domain for which Friedrichs’ inequality is valid); these, in turn, are often needed in stability and domain perturbation analysis (e.g. [20, 28]). Just like, or even more than, in the case of quantum graphs, only in extremely special cases is it possible to determine the eigenvalues exactly.
@@
Theorem 4.2 (Szegő–Weinberger).
@@
The strategy involves symmetrisation (or rearrangement): starting from $\psi_{1}\in H^{1}_{0}(\Omega)$ a positive eigenfunction for $\lambda_{1}(\Omega)$, using tools from geometric measure theory (the coarea formula and the isoperimetric inequality applied to the level sets of $\psi_{1}$) we construct a new function $0<\psi_{1}^{\ast}\in H^{1}_{0}(B)$, a symmetrisation of $\psi_{1}$ such that
@@
4.2 On graphs: Nicaise’ inequality
@@
We wish to understand how these eigenvalues are related to the geometry (and topology, and metric features) of $\Gamma$, akin to the examples sketched in Section 4.1. As discussed in Section 3.4, in general it is essentially impossible to determine the $\mu_{k}$ explicitly (or even to find a convenient representation of the generally extremely complicated transcendental secular equation that characterises them).
@@
Theorem 4.3 (Nicaise’ inequality).
@@
This theorem was originally discovered and proved by Nicaise in 1987 [42], rediscovered by Friedlander in 2005 (with a different proof) [24], and then rediscovered again by Kurasov and Naboko in 2014 [39]. The interpretation in terms of diffusion is that, if the graphs are perfectly insulated, so that no heat can escape, then, among all graphs of given total length, the interval exhibits the slowest convergence (in the usual $L^{2}$-sense) to equilibrium $=$ uniform heat distribution. Actually, there is a corresponding result for discrete graph Laplacians which is even older, going back to the 1970s [23], see also [12, Section 2].
@@
Theorem 4.4 (Nicaise, Dirichlet case).
@@
$\int_{\Gamma}|\psi_{1}^{\prime}|^{2}\,\textrm{d}x\geq\int_{0}^{L}|(\psi_{1}^{\ast})^{\prime}|^{2}\,\textrm{d}x.$
Proof 4.6.
@@
secondly, if $y_{t}\in[0,L]$ is the unique point such that $\psi_{1}^{\ast}(y_{t})=t$, i.e. $\{y_{t}\}=S_{t}^{\ast}=\partial U_{t}^{\ast}$, then the relation $|U_{t}|=|U_{t}^{\ast}|$ implies
@@
where the first line is the coarea formula, the second follows from the first observation, the third from the second observation, and the final line is another application of the coarea formula.
@@
Example 4.8.
If $\Gamma=I=[0,L]$ is just a bounded interval, and $\psi_{k}\sim\lambda_{k}$ is a $k$th eigenfunction of $A_{q}$ on $I$ ($q\in L^{\infty}(I)$), then by Sturm–Liouville theory, $\psi_{k}$ has exactly $k-1$ zeros in the interior of $I$, and thus exactly $k$ nodal domains $\Omega_{1},\ldots,\Omega_{k}$.
@@
$\lambda_{1}(\Omega)=\inf_{\begin{subarray}{c}0\neq u\in H^{1}(\Omega)\\ u(x)=0\,\forall x\in\partial\Omega\end{subarray}}\frac{\int_{\Omega}|u^{\prime}|^{2}+q|u|^{2}\,\textrm{d}x}{\int_{\Omega}|u|^{2}\,\textrm{d}x}.$
@@
Proof 4.11.
By assumption $\psi_{k}$ satisfies the eigenvalue equation $-\psi_{k}=\mu_{k}\psi_{k}$ strongly (indeed pointwise) in $\Gamma$, and hence in $\Omega$. But it also satisfies all vertex conditions in $\Omega$ by construction. Thus, it is equal to some eigenfunction on $\Omega$; in particular, $\mu_{k}=\lambda_{j}(\Omega)$ for some $k\geq 1$.
Now since $\psi_{k}$ does not change sign in $\Omega$, and $\lambda_{1}(\Omega)$ is the only eigenvalue on $\Omega$ with a non-sign-changing eigenfunction, we must have $j=1$.
Proof 4.12 (First proof of Theorem 4.3).
Fix an eigenfunction $\psi_{2}\sim\mu_{2}(\Gamma)$, then $\psi_{2}$ has (at least) two nodal domains $\Omega^{+}$, $\Omega^{-}$; at least one of them, say $\Omega^{+}$, has total length $\leq L/2$. Since $\partial\Omega^{+}\neq\emptyset$, i.e. $\Omega^{+}$ has at least one Dirichlet vertex, by Lemma 4.10 and Theorem 4.4,
@@
Proof 4.14 (Rough idea of proof).
If $\psi_{k}$ has $k$ nodal domains, then at least one has length $\leq L/k$; apply Theorem 4.4. Otherwise, there exists a linear combination of $\psi_{1},\ldots,\psi_{k}$ which does have $k$ nodal domains. Apply the symmetrisation technique to one of these.
Remark 4.15.
@@
On the other hand, the values in Friedlander’s theorem diverge sharply from the values of $\mu_{k}(\Gamma)$ given by Weyl’s law for any fixed graph $\Gamma$. This is starkly at odds with Pólya’s conjecture (see, e.g. [19]) for the Dirichlet and Neumann Laplacian eigenvalues, which posits that the first term in the Weyl asymptotics is always a lower bound for the Dirichlet eigenvalues and an upper bound for the Neumann eigenvalues, on any domain. It seems likely that this is related to the failure of the unique continuation principle mentioned in Section 3.4 and to the nodal domain counts explored in [27].
4.3 Surgery
Second proof of Theorem 4.3: the “doubling trick” (Nicaise/Kurasov–Naboko)
@@ 4.3 Surgery
Duplicate the eigenfunction $\psi_{2}\in H^{1}(\Gamma)$ on each edge to create a new function $\tilde{\psi}_{2}\in\nolinebreak H^{1}(\Gamma_{2})$; then by construction
@@
[…]
Theorem 4.16 (Euler, 1736, “Bridges of Königsberg”).
@@
$\mu_{2}(\Gamma)\geq\mu_{2}(\Gamma_{2})$ $=\inf_{\begin{subarray}{c}0\neq u\in H^{1}(\Gamma_{2})\\ \int_{\Gamma_{2}}u\,\textrm{d}x=0\end{subarray}}\frac{\int_{\Gamma_{2}}|u^{\prime}|^{2}\,\textrm{d}x}{\int_{\Gamma_{2}}|u|^{2}\,\textrm{d}x}=\mu_{2}(\Gamma_{2})$
$\geq\inf_{\begin{subarray}{c}0\neq u\in H^{1}(C)\\ \int_{C}u\,\textrm{d}x=0\end{subarray}}\frac{\int_{C}|u^{\prime}|^{2}\,\textrm{d}x}{\int_{C}|u|^{2}\,\textrm{d}x}=\mu_{2}(C)=\mu_{2}(C).$
@@
Actually, this inequality holds for a more general class of graphs, known as doubly connected graphs (Theorem of Band–Lévy, [8, Theorem 2.1(2)]). Intuitively, graphs which are more connected, i.e. have more non-overlapping paths between any pairs of points, necessarily have larger $\mu_{2}$ and thus faster convergence of diffusion processes to equilibrium. Higher graph connectivity and its effect on the eigenvalues is also explored in [12]. This closely parallels results for discrete graph Laplacians which have been known for many decades [23, 40], where the discrete counterpart of $\mu_{2}$ is even called the algebraic connectivity of the graph.
@@
See also Exercise 3.2 or [13, Section 3]. The process of cutting/gluing is more complicated if other vertex conditions are imposed, since then, unlike with the case of standard conditions, there may be no natural conditions the newly created vertices should satisfy.
Proof 4.20.
@@
(22) now follows from the min-max principle (Theorem 2.4/Remark 3.27).
Corollary 4.21 (Weyl asymptotics).
Let $\Gamma$ be a compact metric graph with total length $L$. Then
@@
Theorem 4.23.
(Cf. [31, Theorem 4.2].) Let $\Gamma$ be a compact metric graph with total length $L>0$ and $m$ edges. Then
@@
Lemma 4.25.
Suppose $\widetilde{\Gamma}$ is formed from $\Gamma$ by deleting a pendant subgraph $\Gamma^{\prime}\subset\Gamma$ (equivalently, $\Gamma$ is formed from $\widetilde{\Gamma}$ by gluing a new graph $\Gamma^{\prime}$ to $\widetilde{\Gamma}$ at a single vertex). Suppose standard conditions are imposed at all vertices of all graphs in question. Then
@@
for all $k\geq 1$.
This lemma can also be proved using a test function argument, see [44, Proposition 3.1]. Its sharpest form can be found in [13, Theorem 3.10]. It is intuitively clear that attaching a pendant subgraph (basically, an extra dead end) should slow the rate of diffusion. The picture is far more complicated if the subgraph being attached or deleted is not a pendant, but rather attached at two or more points, as the following simple example shows.
Example 4.27.
@@
It is left as another exercise to check that the second eigenvalue of an equilateral $m$-flower graph is in fact $\frac{\pi^{2}m^{2}}{L^{2}}$.
4.4 Advanced surgery